following table contains data for a corporations units of production by m of this The year. The two-month forecast is meant to represent an average of the tveh for previous production. The moving average calculation: Units of Month (000) Average Forecast 3,000 2,500 3.500 March Apri May 3.200 2,100 June 3,500 (a) List the two-month average forecast value for January, February, March, April, May, and June. (Use NA when it is not possible to calculate an accurate result.) (b) If there is a more meaningful forecast analysis, detail your recommendations and explain w the results would enhance decision-making. Otherwise, defend the use of the moving average technique in this situation
Answer #1

Two month moving average can be calculated as follows: 4 Units of production Two month Average forecast Month(000) anua Februa March April Ma June 3000 N/A 2500 N/A 3500 3200 2100 3500 10 2750D9+D10)/2 3000 3350 2650 12 13 14 15 16 17 Actual Vs Forecast 19 21 23 24 26 27 29 JanuaryFebruary March April May June _Units of production (000) Two month Averagefor ecast 31 32 It can be seen from the chart that there is lag between the forecasted value and the actual valuee Exponential moving average method gives more weightage to the recent data and hence helps in reducing the lag 35 3637 38 39 In exponential moving average, forecasted value at time t, F(t) in can be written as follows: F(t)-α«T(r-1)+(1-α): F(-1) For 0<a<1 and to-3 Where Y(t-1) is the actual value att-1 and F(t-1) is the forecasted value att-1 For t 1, F(1) Y(1) Given the following data: 41 42 43 45 47 0.2 49 Units of production Month(000) anua Februa March April Ma June Forecasted value 51 52 53 54 3000 3500 3200 2100 3500 3000 3000 SD$48 D51+(1-SD$48)*E51 2900 3020 3056 2864.8 57 Exponential Smothening 59 3500 61 63 JanuaryFebruary March April May _Units of production (000)_Forecasted value 71Formula sheet

A B C D E F G H I J
2
3 Two month moving average can be calculated as follows:
4
5 Ft+1 = (Dt+Dt-1)/2
6
7 a)
8 Month Units of production (000) Two month Average forecast
9 January 3000 N/A
10 February 2500 N/A
11 March 3500 =(D9+D10)/2 =(D9+D10)/2
12 April 3200 =(D10+D11)/2
13 May 2100 =(D11+D12)/2
14 June 3500 =(D12+D13)/2
15
16 b)
17
18
19
20
21
22
23
24
25
26
27
28
29
30
31
32
33 It can be seen from the chart that there is lag between the forecasted value and the actual value.
34 Exponential moving average method gives more weightage to the recent data
35 and hence helps in reducing the lag.
36
37 In exponential moving average, forecasted value at time t, F(t) in can be written as follows:
38
39
40 For 0<α<1 and t>=3
41
42 Where Y(t-1) is the actual value at t-1 and F(t-1) is the forecasted value at t-1
43
44 For t=1, F(1)= Y(1).
45
46 Given the following data:
47
48 α 0.2
49
50 Month Units of production (000) Forecasted value
51 January 3000 =D51
52 February 2500 =$D$48*D51+(1-$D$48)*E51 =$D$48*D51+(1-$D$48)*E51
53 March 3500 =$D$48*D52+(1-$D$48)*E52
54 April 3200 =$D$48*D53+(1-$D$48)*E53
55 May 2100 =$D$48*D54+(1-$D$48)*E54
56 June 3500 =$D$48*D55+(1-$D$48)*E55